OpenStudy (anonymous):

Find at least the first four nonzero terms in a power series expansion about x = 0 for a general solution to the given differential equation: y'-e^x y = 0 ; y(0) = 1

OpenStudy (anonymous):

\[y'-e^x y= 0 ; y(0)=1\] @zepdrix

zepdrix (zepdrix):

Grr I've kind of forgotten how to do these. Let's see... so we want our solution to be of the form:\[\Large\rm y(x)=\sum_{n=0}^{\infty} a_n x^n\]Yes?

OpenStudy (anonymous):

yes, we can either do it explicitly or find the recursive

zepdrix (zepdrix):

So in order to plug stuff in, we also need our derivative.\[\Large\rm y'(x)=\sum_{n=1}^{\infty} n a_n x^{n-1}\]

zepdrix (zepdrix):

explicitly? :o

OpenStudy (anonymous):

i was kidding about that

zepdrix (zepdrix):

So plugging stuff in,\[\Large\rm \sum_{n=1}^{\infty} n a_n x^{n-1}-e^x\sum_{n=0}^{\infty} a_n x^n=0\]Hmm I can't remember this process >.< Do we want the power series expansion of e^x?

OpenStudy (anonymous):

yes I guess so

OpenStudy (anonymous):

the prof hasn't taught us yet, but he wants us to fend for ourselves with this problem.

OpenStudy (anonymous):

i am a babe in the woods right now.

zepdrix (zepdrix):

Example 4 shows it worked out with e^x in the mix. I'm trying to see if I can make sense of it.

OpenStudy (anonymous):

the book jumps around a lot (:((( )

zepdrix (zepdrix):

Oh I thought we'd be doing some type of fancy summation product. But it looks like they're just grabbing the first 4 terms of each expansion and working with those.

zepdrix (zepdrix):

So expanding out e^x as powers of x gives us, \[\Large\rm \sum_{n=1}^{\infty} n a_n x^{n-1}-\left(1+x+\frac{x^2}{2}+\frac{x^3}{6}+\frac{x^4}{24}+...\right) \sum_{n=0}^{\infty} a_n x^n=0\]

zepdrix (zepdrix):

And then we would distribute the summation to the first few terms of our e^x expansion.

zepdrix (zepdrix):

So like for the 1 times summation we would get,\[\Large\rm a_o+a_1 x+a_2 x^2+a_3 x^3+...\]For the x times summation we would get,\[\Large\rm a_o x+a_1 x^2+a_2 x^3+a_3 x^4+...\]For the x^2/2 times summation we would get,\[\Large\rm a_o \frac{x^2}{2}+a_1 \frac{x^3}{2}+a_2 \frac{x^4}{2}+a_3 \frac{x^5}{2}+...\]

zepdrix (zepdrix):

Do that up to the 4th degree term and also expand out the derivative summation as well. Then group like terms together and find your recurrence relation. It's going to be long and tedious :O It prolly won't be as bad on paper. Typing it all out would be a nightmare though lol